A yardstick measures 0.5 by 2 by 36 inches. How many yardsticks will fit in a box 3.5 inches wide and 36 inches high, if the length of the box is 18 inches?
Answer by checkley77(12844)

.5*2=1 in^2 is the end area of the yardstick.
3.5*18=63 in^2 is the end area of the 36 inch box.
63/1=63 yardsticks can fit in this box.